Make New Entry,
Make Followup Entry
User name V. Sulkosky
Log entry time 17:27:23 on October 13, 2009
Entry number 294756
This entry is a followup to: 294575
keyword=Run Coordinator summary: October 12 to 13 (day shift)
RC report (Oct. 12 Owl shift ~ Oct. 13 Day shift)
Monday, Oct. 12
- 00:00 Beam off due to IOCHLA needing to be rebooted in Hall A.
- Controlled access to reboot IOCHLA and change Compton FADC
- 00:50 Beam restored and checking raster.
- 00:57 Resume 100 uA production with trippy beam.
- 01:00-05:30 Compton background rate investigative work.
- 05:30 Compton rate work concludes without success.
- 06:51 Beam IHWP is inserted and began slug # 22.
- 09:10 Beam off for spot move and injector work for Compton rates.
- 17:40 Injector work completed and Compton background rates are
more reasonable at 200-300 Hz/uA.
- 18:37 After beam restoration and beam profile checks, production
resumes at 100 uA with Compton production.
- The rest of swing shift was very smooth with only 14
beam trips between 19:00 and 24:00.
Tuesday, Oct. 13, (Owl and Day)
- Production running with fairly stable beam and okay Compton rates.
- 00:013-02:23 Compton tests runs with electron detector.
- 02:23 Resume Compton production running.
- Fairly smooth production running into Day shift.
- 11:30 Compton rates started to increase and brought back down to
200 Hz/uA by adjusting MBH2I03V.
- 13:09 Begin setting up for Moller measurement.
- 13:37-14:09 Controlled access to reboot Moller crate.
- Day shift ends with Moller in progress.
Coulombs So Far
- We are currently on slug # 22 (IHWP IN)
- Coulombs taken in this time period (since Sunday 09:00 am): 7.5 C
- Coulombs taken in all so far: 129.4 C as of Tuesday 4:00 am.
- Please always check online hists and compare them to the
- If you find something strange or wrong, please HALOG it so
that the expert can check it.
- Target Operator, please always check target PID system and
strip plots to make sure all the systems work.
The cryo flow rate around 25.2g/s and JT value is around
55.5 since this is better for Lumi widths.
The flow rate and JT value setting are sensitive to the outside
temperature. Hence the
settings may change from morning to evening.
Please be patient when you adjust the JT value since it
may take a few minutes to make the cryo flow rate stable.